Circular No. 7291

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

X-RAY TRANSIENT IN M31 S. S. Murray, M. R. Garcia, and F. A. Primini, Harvard-Smithsonian

Center for Astrophysics, report: “We have discovered a bright new x-ray transient located 13′′ west of the radio nucleus of M31. A 15 000-s observation with the Chandra ACIS-I detector, centered on Oct. 13.23 UT, shows a new source at α = 0^{h}42^{m}42^{s}.18, δ = +41^{o}16′08′′.4 (equinox 2000.0; 90percent error radius 2′′.0), with a luminosity of ∼ 10^{38} erg/s in the Chandra bandpass 0.5–10 keV. This position is derived from a plate solution using eight globular clusters detected in the Chandra image and positions referenced to the HST Guide Star Catalog (Magnier et al. 1993, A.Ap. 272, 695). The last ROSAT HRI observation of this region in Jan. 1996 shows no source at this location, to an upper limit of 5 × 10^{35} erg/s. Observations at other wavelengths are encouraged in order search for a possible counterpart.”

XTE J1859+226 D. Dal Fiume, Istituto TeSRE, CNR Bologna, communicates (together

with F. Frontera, M. Orlandini, L. Amati, S. Del Sordo, L. Stella, T. Belloni, D. Ricci, M. Capalbi, and M. R. Daniele, on behalf of a large collaboration): “BeppoSAX observed the transient source XTE J1859+226 on Oct. 15.086– 15.405 UT for ∼ 14 000 s in the MECS and HPGSPC and for 7000 s in the PDS. The source was clearly detected up to 200 keV. The 2–200keV spectrum is roughly approximated by an absorbed power law with photon index of ∼ −2.8 and N_{H} ∼ 8 × 10^{21} cm^{−2}. This spectral slope is consistent with that measured with BATSE (IAUC 7282) and steeper than that measured with RXTE (IAUC 7274). The source flux was 1.2 × 10^{−8} erg s^{−1} cm^{−2} (560 mCrab) in the energy band 2–10 keV, and was 3.3 × 10^{−9} erg s^{−1} cm^{−2} (190 mCrab) in the energy band 20–100 keV. A power-spectral-density estimate shows a prominent quasiperiodic oscillation (QPO) centered at 5.49 ± 0.04 Hz (FWHM = 0.7 ± 0.1 Hz) both in the MECS and in the HPGSPC data, with a fractional r.m.s. amplitude of ∼ 5 percent in the band 1.5–10 keV. A less-prominent QPO peak is observed near ∼ 0.6 Hz only in the HPGSPC data, while we do not find evidence for the 0.45-Hz QPO observed by RXTE (IAUC 7274). These timing and spectral characteristics are reminiscent of a blackhole candidate in outburst. Additional BeppoSAX observations are planned.”

1999 October 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green